# Tackling the Accuracy-Interpretability Trade-off in a Hierarchy of Machine Learning Models for the Prediction of Extreme Heatwaves

### v1

@@ 6 Discussion
By testing increasingly complex architectures, we were able to quantify the extent to which higher complexity leads to a better performance. As pointed out in Rudin (2019), the benefits of complexity are often overrated, and indeed, when data is scarce, the best performance was achieved by the simple linear regression of the Gaussian approximation.
@@
This simple model remained competitive with the more complex ones, even with much larger amounts of data. Moreover, when the Scattering Network uses only the coarse-grained input, it becomes itself a simple linear regression, and it has the same skill as the Gaussian approximation while using 64 times less scalar features. This highlights the potential for even further simplification of the regression task without loss of performance from the GA baseline.
@@
For the task of forecasting heatwaves, the slightly increased complexity of the Intrinsically interpretable Neural Network did not give any benefits, neither from the point of view of performance nor of additional insight into the physical processes. However, the design of the architecture is very interesting, reimagining a simple fully connected neural network with a bottleneck in terms of optimal linear projection of the data. It is then easy to relax the bottleneck from a single optimal index to a set of optimal indices. This could be particularly useful when studying problems where the effective dynamics can be well expressed non-linearly in a low-dimensional space. A possible example could be the study of the Madden-Julian Oscillation (Delaunay and Christensen, 2022). Compared to standard principal component analysis, the IINN would give linear projections that are tailored to the regression task, rather than the ones that best explain the variance of the input data.
We did try to train IINN architectures with more than one projection pattern on the heatwave prediction task, but the results were not better than the ones presented in this work.
@@
Indeed, going beyond the skill of the linear model requires several centuries of data to properly train more flexible architectures, like Convolutional Neural Networks. With post-hoc explainability methods, we found that the CNN essentially refines and modulates the projection pattern of the linear model. However, we advocate for caution in using these methods. If we had just looked at the CNN and its post-hoc explanations, we would have been happy to see them highlighting dry soil, the anticyclone over France and activity over the Atlantic. Only by having a hierarchy of models we realize that these easily recognizable features were already present in the linear model, and thus the extra skill that we gained lies in second order details. Pinpointing such details is not easy, and we could not draw robust conclusions. Nevertheless, the explainability maps do provide hints of where extra sources of predictability might be, which can be useful for future investigations with more rigorous methods.
The architecture of the ScatNet proposed in this work allows to overcome the limitations of post-hoc tools, providing global interpretability like GA and IINN. Importantly, the ScatNet has the same skill as the CNN, so we can say that there is nothing significant in the predictions of the CNN that the ScatNet misses.
@@
Our analysis shows that the linear model contributes to 65% of the prediction of the ScatNet, with the remaining information being encoded in oscillations in the geopotential height anomaly field at sub-synoptic scales. Even more interestingly, we are able to visualize the individual contribution of oscillations in different directions and geographical locations. The precise physical interpretation of these oscillations goes beyond the scope of this paper, but we feel this is a very promising direction for the discovery of new physics.
On a more technical note, in our work the Scattering Transform is applied to the geopotential height field by treating it as a two-dimensional image with pixels of constant size in latitude and longitude. A natural improvement is then to use the Scattering Transform on a sphere (McEwen et al., 2021) to better account for Earth’s geometry. Then, ScatNet could be integrated as an interpretable module (Yang et al., 2024) into more complex models, serving as a preprocessing step to give a sparse representation of a geophysical field. For instance, it could potentially act as the initial layer in an encoder-decoder architecture. Furthermore, the mean scattering spectrum could be used as a penalization term in loss functions for deep learning weather and climate emulators to promote physically consistent forecasts.
Finally, a general remark is that in this work we approximated, across all models, the conditional distribution of the heatwave amplitude as a Gaussian distribution, of which we estimate the mean and the standard deviation. This may be a factor limiting performance in highly non-linear problems where the conditional distributions are potentially skewed or even multimodal. A possible fix could be to use a non-parametric approach such as quantile regression (Zhang et al., 2018) or other types of parametric approaches, for instance relying on extreme value theory (Cisneros et al., 2024). However, this has the effect of complicating network architectures and thus hindering interpretability, which was the main focus of this work.

### v2

@@ 6 Discussion
By testing increasingly complex architectures, we were able to quantify the extent to which higher complexity leads to a better performance. As pointed out in Rudin (2019), the benefits of complexity are sometimes overrated, and indeed, when data is scarce, the best performance was achieved by the simple linear regression of the Gaussian approximation. This simple model remained competitive with the more complex ones, even with much larger amounts of data. Moreover, when the Scattering Network uses only the coarse-grained input, it becomes itself a simple linear regression, and it has the same skill as the Gaussian approximation while using 64 times fewer scalar features. This highlights the potential for even further simplification of the regression task without loss of performance from the GA baseline.
@@
For the task of forecasting heatwaves, the slightly increased complexity of the Intrinsically interpretable Neural Network did not give any benefits, neither from the point of view of performance nor of additional insight into the physical processes. However, the design of the architecture is very interesting, reimagining a simple fully connected neural network with a bottleneck in terms of optimal linear projection of the data. It is then easy to relax the bottleneck from a single optimal index to a set of optimal indices. This could be particularly useful when studying problems where the effective dynamics can be well expressed non-linearly in a low-dimensional space. A possible example could be the study of the Madden-Julian Oscillation (Delaunay and Christensen, 2022). Compared to standard principal component analysis, the IINN would give linear projections that are tailored to the regression task, rather than the ones that best explain the variance of the input data. We did try to train IINN architectures with more than one projection pattern on the heatwave prediction task, but the results were not better than the ones presented in this work.
@@
Going beyond the skill of the linear model requires several centuries of data to properly train more flexible architectures, like Convolutional Neural Networks. With post-hoc explainability methods, we found that the CNN essentially refines and modulates the projection pattern of the linear model. However, we advocate for caution in using these methods. If we had just looked at the CNN and its post-hoc explanations, we would have been happy to see them highlighting dry soil, the anticyclone over France and activity over the Atlantic. Only by having a hierarchy of models we realize that these easily recognizable features were already present in the linear model, and thus the extra skill that we gained lies in second order details. Pinpointing such details is not easy, and we could not draw robust conclusions. Nevertheless, the explainability maps do provide hints of where extra sources of predictability might be, which can be useful for future investigations with more rigorous methods.
@@
The architecture of the ScatNet proposed in this work allows to overcome the limitations of post-hoc tools, providing global interpretability like GA and IINN. Importantly, the ScatNet has the same skill as the CNN, so we can say that there is nothing significant in the predictions of the CNN that the ScatNet misses. Our analysis shows that the linear model contributes to 65% of the prediction of the ScatNet, with the remaining information being encoded in oscillations in the geopotential height anomaly field at sub-synoptic scales. Even more interestingly, we are able to visualize the individual contribution of oscillations in different directions and geographical locations. The precise physical interpretation of these oscillations goes beyond the scope of this paper, but we feel this is a very promising direction for the discovery of new physics.
As a technical note, in this work we apply the Scattering Transform to the geopotential height field, treating it as a 2D image. However, a spherical adaptation (McEwen et al., 2021) would better account for Earth’s geometry. ScatNet could then serve as an interpretable preprocessing step, providing a sparse representation of geophysical fields—either as an encoder’s initial layer or as a regularization term to enforce physically consistent forecasts. Meanwhile, our Gaussian approximation of the conditional distribution of heatwave amplitude, , may limit performance in nonlinear cases where distributions are skewed or multimodal. Other approaches like quantile regression (Zhang et al., 2018) or extreme value theory (Cisneros et al., 2024) could improve flexibility but at the cost of increased model complexity and reduced interpretability, which was our primary focus.
From a more physical standpoint, in this paper we call heatwaves any points for which , which naturally means our heatwave samples are correlated, and often belong to a longer-lasting or more instense heatwave. This means we are more sensitive to stationary features, characteristic of the continuation of a heatwave rather than its onset. Similarly, as for any statistical model, finding a prediction that works well for all the stages of a heatwave tends to average out the dynamics. Ideally one could train different models on the different stages, but that comes with the issue of precise definitions of said stages and reduced dataset size. We tried a-posteriori to look at the EGFI maps of the CNN specifically at heatwave onsets. If onset dynamics are very different from continuation, the CNN, being highly non-linear, should be able to capture them both. However, there was no obvious difference with respect to EGFI maps captured at the center of long lasting heatwaves.
This points out that explainable AI methods are only the first step in understanding the problem at hand. Once interesting features are identified, we should complement the analysis by computing composite statistics, conditioned on the occurrence of those features rather than the target event itself. Alternatively, we could perform denial experiments, where those features are either the only information a network can see or are instead masked out. This would allow to precisely and reliably quantify their effective importance for prediction. Another very interesting possibility would be to run idealized experiments with climate models, where, for instance, we explicitly force the sub-synoptic oscillations found by the ScatNet. This would require first some analysis work to identify the conditions which favors those oscillations. These experiments are beyond the scope of this paper, which was to investigate the tradeoff between performance and interpretability. However, we think the results presented in this work will be useful to improve, in future studies, their interpretation.
@@ 7 Conclusions
However, these conclusions hold for this specific problem, where nonlinearity is moderate; for more nonlinear tasks, black-box models like CNNs could be more advantageous.
@@ Acknowledgements.
- DelSole and Banerjee (2017) DelSole, T., and A. Banerjee, 2017: Statistical Seasonal Prediction Based on Regularized Regression. Journal of Climate, 30 (4), 1345–1361, 10.1175/JCLI-D-16-0249.1.
@@
- Materia et al. (2024) Materia, S., and Coauthors, 2024: Artificial intelligence for climate prediction of extremes: State of the art, challenges, and future perspectives. WIREs Climate Change, 15 (6), e914, 10.1002/wcc.914.
@@
- Stevens et al. (2021) Stevens, A., R. Willett, A. Mamalakis, E. Foufoula-Georgiou, A. Tejedor, J. T. Randerson, P. Smyth, and S. Wright, 2021: Graph-Guided Regularized Regression of Pacific Ocean Climate Variables to Increase Predictive Skill of Southwestern U.S. Winter Precipitation. Journal of Climate, 34 (2), 737–754, 10.1175/JCLI-D-20-0079.1.
@@
- Zschenderlein et al. (2019) Zschenderlein, P., A. H. Fink, S. Pfahl, and H. Wernli, 2019: Processes determining heat waves across different European climates. Quarterly Journal of the Royal Meteorological Society, 145 (724), 2973–2989, 10.1002/qj.3599.